OpenStudy (anonymous):

Solve the equation: 87+2x/216+3x=.5

OpenStudy (saifoo.khan):

please use brackets, i can solve.

OpenStudy (anonymous):

\[87+2x \div216+3x =.5\]

OpenStudy (saifoo.khan):

Hi Anyways! :)

OpenStudy (anonymous):

hi hi :)

OpenStudy (saifoo.khan):

325x 0.5 - 87 ---- = 108 325x = 9342 x = 9342/325 x = 28.74

OpenStudy (anonymous):

Thank you saifoo :)Thank you saifoo :)

OpenStudy (saifoo.khan):

uw.